Let CTO,1] 3.(12 pts) Let f=f(x)-mx-1 and g = g have the inner product (x)-4x + m be two functions in q0,1]. Find the exact value(s) of m for which f and g are orthogonal.
Answer #1

Solution:

\small f=f\left ( x \right )=mx-1

\small g=g\left ( x \right )=4x+m

\small f and  \small g are orthogonal

\small \therefore \left \langle f,g \right \rangle=0

\small \therefore \int_{0}^{1}f\left ( x \right )g\left ( x \right )dx=0

\small \therefore \int_{0}^{1}\left ( mx-1 \right )\left ( 4x+m \right )dx=0

\small \therefore \int_{0}^{1}\left ( 4mx^{2}+m^{2}x-4x-m \right )dx=0

\small \therefore \left [ 4m\cdot \frac{x^{3}}{3}+m^{2}\cdot \frac{x^{2}}{2}-4\cdot \frac{x^{2}}{2}-mx\right ]_{0}^{1}=0

\small \therefore \frac{4m}{3}+\frac{m^{2}}{2}-2-m=0

\small \therefore \frac{4m}{3}+\frac{m^{2}}{2}-2-m=0

\small \therefore 3m^{2}+2m-12=0

\small \therefore m=\frac{\sqrt{37}-1}{3}\: ,\: \frac{-\sqrt{37}-1}{3}.